# A Carbon Tracking Model for Federated Learning: Impact of Quantization and Sparsification thanks: This paper is funded by the EU in the call HORIZON-HLTH-2022-STAYHLTH-01-two-stage under grant agreement No 101080564

### v1

@@ II-A Energy tracking for Parameter Server based FL (FA)
where $\mathrm{C}^{(\text{FA})}_{k}(t_{i})=E_{k,i}^{\text{(FA)}}\cdot I_{k}+\mathrm{C}_{k}^{(\text{FA})}(t_{i-1})$ is the carbon footprint of device $k$ at round $i$ while $\mathrm{C}_{0}^{(\text{PS})}(t_{i})=E_{0,i}^{\text{(PS)}}\cdot I_{0}+\mathrm{C}_{0}^{(\text{PS})}(t_{i-1})$ is the carbon footprint of the PS at round $i$.
@@ II-B Energy tracking for decentralized CFA
The total carbon consumption under the CFA scheme can be quantified as
@@
with $C_{k}^{(\text{CFA})}(t_{i})=E_{k,i}^{(\text{CFA})}\cdot I_{k}+C_{k}^{(\text{CFA})}(t_{i-1})$ being the carbon footprint of device $k$ at round $i$.
@@ III Quantization and parameter selection in FL
where $\mathcal{Q}(.)$ denotes the overall compression function, while $f_{s}(.)$ and $f_{q}(.)$ indicate the operators required for top-$t$ sparsification and probabilistic quantization, respectively. The sparsification operator $f_{s}(.)$ outputs a new representation $\widetilde{\mathbf{w}}=f_{s}(\mathbf{w})$ that selects the $t$ largest absolute values of $\mathbf{w}$, and sets all other entries to 00. Quantization $f_{q}(.)$ encodes the sparsification outputs $\widetilde{\mathbf{w}}$ via a randomized rounding operation. By setting the output quantization bits to $N_{b}\leq N_{bc}$, with (typical) $N_{bc}=32$ bits, the element $\bar{w}_{n}$ of $\bar{\mathbf{w}}$ is defined as [22]
@@
where $\delta=\tfrac{t}{N_{P}}$ represents the fraction of the model parameters selected by top-$t$ sparsification function $f_{s}(\mathbf{v})$, while $\tfrac{N_{b}}{N_{bc}}$ sets the quantization level according to $f_{q}(.)$ for each parameter.
@@ IV Carbon footprint evaluation
This section discusses the main factors that are expected to steer the choice between vanilla FA and CFA paradigms towards sustainable designs. The goal is to provide a first look into the impact of quantization and sparsification of model parameters on carbon emissions and learning accuracy, for varying communication energy efficiencies typically found in wireless communication systems. We consider two scenarios with increasing model size and training dataset complexities. The first case (Sec. IV-A) focuses on the MNIST classification task [8], while the second (Sec. IV-B) concentrates on the CIFAR10 learning problem [9]. Energy and carbon footprints are influenced by the PS and device hardware configurations. Table I summarizes the main parameters used by the carbon tracking tool under the two scenarios. For the PS (FA), we used a CPU (Intel i7 8700K, $3.7$ GHz, $64$ GB, GPU not used). A realistic pool of $K=10$ resource-constrained FL learners is adopted, namely Jetson Nano boards based on a low-power CPU (ARM-Cortex-A57 and GPU 128-core Maxwell). Each device is designed to track its carbon emissions independently as a function of the estimated CI $I_{k,i}$ which is assumed as constant for all FL rounds.
@@ IV-A Impact of sparsification, energy and carbon efficiencies
Fig. 2 reports the validation accuracy obtained by the CFA and FA schemes under the two carbon budgets. For stringent emission constraints, i.e., $C_{\text{tot}}=10$ gCO2-eq, CFA provides better performances when $\delta=10\%-50\%$ and $N_{b}=8$ bits, while for $C_{\text{tot}}=100$ gCO2-eq FA achieves higher accuracy regardless of the number of parameters shared and the quantization bits employed. ML model compression is more critical in CFA rather than in FA since it reduces the sidelink channel use. FA is not that much affected by the specific choice of the compression parameters as long as the carbon budget is high enough, i.e., $C_{\text{tot}}=100$ gCO2-eq. On the other hand, CFA schemes must generally employ a more aggressive compression, i.e., $\delta=10\%$ and $N_{b}=8$ bits, to guarantee a reasonable target accuracy. For FA, a good choice for the compression parameters is $\delta=10\%$ and $N_{b}=16$ bits as achieving the highest accuracy under all carbon budgets.
@@
Fig. 3 analyzes the impact of the communication energy efficiency as well as the CI on the learning accuracy under the same carbon budget $\mathrm{C}_{\text{tot}}=10$ gCO2-eq. The energy efficiency $\text{EE}_{\text{COM}}$ varies from $5$ Kbit/Joule up to $50$ Kbit/Joule, with the last case corresponding to 5G micro-cell delivery and WiFi. The results consider $\delta=10\%$ and $N_{b}=16$ bits for FA, while $\delta=10\%$ and $N_{b}=8$ bits for CFA following the previous analysis. In the same figure, we also study how the PS location and its CI $I_{0}$ affect the carbon consumption. Three CI values are considered, namely $I_{0}=\{0.193,0.449,0.952\}$ KgCO2-eq/kWh, modeling different energy grid efficiencies. These values correspond to the geographical regions of Finland, Italy, and Poland, respectively. Comparing the results, CFA is shown to outperform FA for all CI terms, provided that the energy efficiency is below $25$ Kbit/Joule. On the other hand, when the communication is more efficient, centralized FA should be preferred. High carbon intensities, i.e., $I_{0}=0.952$ KgCO2-eq/kWh, make FA tools more susceptible to accuracy losses due to the high energy costs. Interestingly, even when the PS is located in the same region of the devices, CFA strategies are to be preferred as more accurate compared PS-based solutions.

### v2

@@ II-A Energy tracking for Parameter Server based FL (FA)
where $\mathrm{C}^{(\text{FA})}_{k}(t_{i})=E_{k,i}^{\text{(FA)}}\cdot I_{k,i}+\mathrm{C}_{k}^{(\text{FA})}(t_{i-1})$ is the carbon footprint of device $k$ at round $i$ while $\mathrm{C}_{0}^{(\text{PS})}(t_{i})=E_{0,i}^{\text{(PS)}}\cdot I_{0,i}+\mathrm{C}_{0}^{(\text{PS})}(t_{i-1})$ is the carbon footprint of the PS at round $i$.
@@ II-B Energy tracking for decentralized CFA
The total carbon consumption under CFA can be quantified as
@@
with $C_{k}^{(\text{CFA})}(t_{i})=E_{k,i}^{(\text{CFA})}\cdot I_{k,i}+C_{k}^{(\text{CFA})}(t_{i-1})$ being the carbon footprint of device $k$ at round $i$.
@@ III Quantization and parameter selection in FL
where $\mathcal{Q}(.)$ denotes the overall compression function, while $f_{s}(.)$ and $f_{q}(.)$ indicate the operators required for top-$t$ sparsification and probabilistic quantization, respectively. The sparsification operator $f_{s}(.)$ outputs a new representation $\widetilde{\mathbf{w}}=f_{s}(\mathbf{w})$ that selects the $t$ largest absolute values of $\mathbf{w}$, and sets all other entries to 00. Quantization $f_{q}(.)$ encodes the sparsification output $\widetilde{\mathbf{w}}$ via a randomized rounding operation. By setting the output quantization bits to $N_{b}\leq N_{bc}$, with (typical) $N_{bc}=32$ bits, the element $\bar{w}_{n}$ of $\bar{\mathbf{w}}$ is defined as [22]
@@
where $\delta=\tfrac{t}{N_{P}}$ represents the fraction of the model parameters selected by top-$t$ sparsification function $f_{s}(.)$, while $\tfrac{N_{b}}{N_{bc}}$ sets the quantization level according to $f_{q}(.)$ for each parameter.
@@ IV Carbon footprint evaluation
This section discusses the main factors that are expected to steer the choice between vanilla FA and CFA paradigms towards sustainable designs. The goal is to provide a first look into the impact of quantization and sparsification of model parameters on carbon emissions and learning accuracy, for varying communication energy efficiencies typically found in wireless communication systems. We consider two scenarios with increasing model size and training dataset complexities. The first case (Sec. IV-A) focuses on the MNIST classification task [8], while the second (Sec. IV-B) concentrates on the CIFAR10 learning problem [9]. Energy and carbon footprints are influenced by the PS and device hardware configurations. Table I summarizes the main parameters used by the carbon tracking tool under the two scenarios. For the PS (FA), we used a CPU (Intel i7 8700K, $3.7$ GHz, $64$ GB, GPU not used). A realistic pool of $K=10$ resource-constrained FL learners is adopted, namely Jetson Nano boards based on a low-power CPU (ARM-Cortex-A57 and GPU 128-core Maxwell). Their energy expenditure parameters when implementing the FA and CFA schemes reviewed in Sec. III are summarized in Table I. Each device is designed to track its carbon emissions independently as a function of the estimated CI $I_{k,i}=I_{k}$ which is assumed as constant for all FL rounds.
@@ IV-A Impact of sparsification, energy and carbon efficiencies
Fig. 2 reports the validation accuracy obtained by the CFA and FA schemes under the two carbon budgets. For stringent emission constraints, i.e., $C_{\text{tot}}=10$ gCO2-eq, CFA provides better performances when $\delta=\{10\%,50\%\}$ and $N_{b}=8$ bits, while for $C_{\text{tot}}=100$ gCO2-eq FA achieves higher accuracy regardless of the number of parameters shared and the quantization bits employed. ML model compression is more critical in CFA rather than in FA since it reduces the sidelink channel use. FA is not that much affected by the specific choice of the compression parameters as long as the carbon budget is high enough, i.e., $C_{\text{tot}}=100$ gCO2-eq. On the other hand, CFA schemes must generally employ a more aggressive compression, i.e., $\delta=10\%$ and $N_{b}=8$ bits, to guarantee a reasonable target accuracy. For FA, a good choice for the compression parameters is $\delta=10\%$ and $N_{b}=16$ bits as achieving the highest accuracy under all carbon budgets.
@@
Fig. 3 analyzes the impact of the communication energy efficiency as well as the CI on the learning accuracy under the same carbon budget $\mathrm{C}_{\text{tot}}=10$ gCO2-eq. The energy efficiency $\text{EE}_{\text{COM}}$ varies from $5$ Kbit/Joule up to $50$ Kbit/Joule, with the last case corresponding to 5G micro-cell delivery and WiFi. The results consider $\delta=10\%$ and $N_{b}=16$ bits for FA, while $\delta=10\%$ and $N_{b}=8$ bits for CFA following the previous analysis. In the same figure, we also study how the PS location and its CI $I_{0}$ affect the carbon consumption. Three CI values are considered, namely $I_{0}=\{0.193,0.449,0.952\}$ KgCO2-eq/kWh, modeling different energy grid efficiencies. These values correspond to the geographical regions of Finland, Italy, and Poland, respectively. Comparing the results, CFA is shown to outperform FA for all CI terms, provided that the energy efficiency is below $25$ Kbit/Joule. On the other hand, when the communication is more efficient, centralized FA should be preferred. High carbon intensities, i.e., $I_{0}=0.952$ KgCO2-eq/kWh, make FA tools more susceptible to accuracy losses due to the high energy costs. Interestingly, even when the PS is located in the same region of the devices, CFA strategies are to be preferred as more accurate compared to PS-based solutions.